# Genetic algorithm and edge-colorings of complete graphs with connected classes

### v2

@@ Abstract
In this study, the Rank Genetic Algorithm (Rank GA) was adapted to address a problem in the field of Chromatic Graph Theory: the connected pseudo-pseudoachromatic index. We successfully improved several previously known lower bounds of that index. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. The findings highlight the interdisciplinary impact of the use of genetic algorithms in theoretical mathematical research, particularly in the solution of complex combinatorial problems. Our work demonstrates that computational methods can solve problems that are intractable to humans, leading to scientific advances in many areas. Moreover, the problem we addressed showcases the robustness of the Rank GA in finding optimal solutions within a reasonable time frame, as it consistently avoids getting stuck in local optima.
Keywords: Genetic Algorithm, Rank GA, Chromatic Graph Theory, Edge-Coloring, Connected Classes
@@ 1 Introduction
On the other hand, within the field of mathematics, particularly Graph Theory, we focus on graph coloring research, exploring the specialized field of Chromatic Graph Theory. Our specific attention is directed towards the connected-pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$. Determining $\psi_{c}^{\prime}(K_{n})$ is difficult, and so far, there have been very few exact results for $\psi_{c}^{\prime}(K_{n})$ of complete graphs with connected classes. However, it has been possible to bound it using analytical methods [2].
In this interdisciplinary study, our objective was to refine the known bounds of $\psi_{c}^{\prime}(K_{n})$ exploiting the capabilities of the Rank GA. We successfully improved several of the previously known lower bounds. To validate the effectiveness of these new bounds, we performed an analytical approximation, confirming their validity. This synergy between Genetic Algorithms and mathematical theory contributes to the intersection of practical problem-solving and theoretical advancements.
The structure of this paper is as follows. In Section 2, we provide an overview of previous work in which a GA has been successfully used to solve graph theory problems. In Section 3, we state the problem and include the necessary definitions to understand it. In Section 4, we describe the Rank GA and in Section 5, we present the obtained results. Finally, Section 7 contains the conclusions.
2 Background
@@
In recent decades, GAs have emerged as effective tools to address problems of graph theory with practical applications. Specifically, various GAs have been developed to tackle the Graph Coloring Problem (GCP), a fundamental issue in graph theory. In the simplest instance of the GCP, the objective is to assign colors to the vertices of a graph in such a way that no two adjacent vertices share the same color. The goal is to find a valid coloring of the graph with the minimal number of colors. Given its NP-hard nature and numerous practical applications, the GCP stands out as the most extensively researched combinatorial optimization problem in both computer science and mathematics [7]. In [12], you can find the most significant work in which GAs have been used to address the GCP. In addition to GCP, GAs have also been applied to other graph problems, among the most recent works are: the Shortest Path Problem [6], the Traveling Salesman Problem [13], Maximum Matching in bipartite graphs [11], and the Maximum Edge-Disjoint Path Problem [9].
@@
Although GAs have demonstrated effectiveness in the area of graph theory for solving practical problems, to the best of our knowledge, their utilization for theoretical problems is not as widespread. We are aware of the following studies where a genetic algorithm played a crucial role in addressing theoretical challenges: In [5] the Rank GA was used to find rainbow $t$-colorings of the family of Moore cages with girth 6: $(k;6)$, achieving an improvement in the upper bound for the 4-rainbow coloring of a $(4;6)$-cage. In addition, in [8], the Rank GA was adapted with recombination and mutation operators from graph theory. These operators enabled the discovery of Hajós operations to construct the symmetric 5-cycle from the complete symmetric digraph of order 3 using only Hajós operations. Finally, in [4] real solutions were identified by a genetic algorithm, thus contributing to the validation of a conjecture. These studies illustrate that GAs can make valuable contributions to the field of mathematics. This, in turn, inspired us to tackle the challenge of enhancing bounds in edge-colorings of complete graphs with connected classes.
@@ 3 Problem Statement
The maximum number of colors in a connected complete coloring of a complete graph $K_{n}$ of order $n$ that can be achieved is called the connected pseudo-achromatic index and is denoted $\psi_{c}^{\prime}(K_{n})$. There are certain values of $n$ for which it is known, and for others, only the lower and upper bounds are known [1, 2]. The problem we address here, in particular, consists in improving the known bounds [2] for some values of $n$.
@@ 3.1 Edge-Colorings of a Complete Graph
The connected-pseudoachromatic index $\psi^{\prime}_{c}(G)$ of a graph $G$ is the maximum number of colors $k$ for which a connected complete edge-coloring exists in the graph $G$ using $k$ colors, when $G$ is connected; otherwise, it is defined as the maximum index over its connected components. When we color vertices and each chromatic class is connected, the parameter is also known as the Hadwiger number of $G$.
@@
In [2] it is shown that $\psi_{c}(n)$ can be bounded by analytical methods obtaining that $\psi_{c}(n)=\theta(n^{3/2})$. Table 1 shows the known bounds of $\psi_{c}(n)$, for complete connected graphs of $2\leq n\leq 31$ vertices. With the use of the computer, complete colorings can be achieved, and thus improve the lower bound of $\psi_{c}(n)$.
@@ 4.1 The Genetic Operators of the Rank GA
Once individuals are ranked based on their fitness, from best to worst, the Rank GA applies the following operations iteratively until a stopping criterion is met:
  • Rank-based selection
  • Sort and rank
  • Rank-based recombination
  • Evaluate, sort, and rank
  • Rank-based mutation
  • Evaluate, sort, and rank
@@ 5 Results
Table 2 presents the improved lower bounds obtained through the Rank GA for various values of $n$. The known upper bounds for $\psi_{c}^{\prime}(K_{n})$ are also shown for reference in Table 2. By comparing the improved lower bounds with the existing lower and upper bounds, we observed significant reductions in the gap between them.
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper $1$ $3$ $4$ $6$ $7$ $10$ $14$ $18$ $22$ $25$ Lower Improved $13$ $16$ $18$ Lower $1$ $3$ $4$ $6$ $7$ $10$ $11$ $12$ $13$ $14$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Upper $28$ $31$ $34$ $37$ $40$ $45$ $51$ $57$ $63$ $70$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Lower $19$ $26$ $27$ $28$ $29$ $30$ $31$ $32$ $33$ $42$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Upper $74$ $78$ $82$ $86$ $90$ $94$ $98$ $102$ $106$ $116$ Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Lower $43$ $44$ $45$ $46$ $47$ $48$ $49$ $50$ $51$ $93$
Table 2: Improved lower bounds with the Rank GA for $\psi_{c}(n)$.
@@ 6 On the Upper Bound
A natural question is: Are the bounds obtained by the Rank GA good? For this, we give an approximation of the upper bound of the $\psi_{c}(n)$ function by doing an analysis of the average degree with which two acyclic chromatic classes are incident.
Experimentally, only some chromatic classes contain cycles. Table 3 shows the average degrees of incidence between chromatic classes for $n\leq 30$. To achieve a range of heuristically acceptable values, we obtain an approximation of $\psi_{c}(n)$ under the hypothesis that each chromatic class contains at least one cycle, the average degree of which is at least 2.
$n$ num colors average degrees of incidence
8 11 1.563636
9 13 1.576923
10 16 1.391666
11 18 1.398692
12 21 1.361904
13 26 1.120000
14 27 1.333333
15 28 1.550264
16 29 1.894088
17 32 1.616935
18 35 1.591596
19 37 1.696696
20 39 1.716599
21 43 1.660022
22 46 1.662801
23 50 1.651428
24 52 1.718702
25 54 1.758909
26 55 1.806060
27 59 1.832846
28 62 1.875198
29 64 1.875000
30 66 1.920279
Table 3: The average degrees of incidence between the chromatic classes obtained by the Rank GA.
@@
Furthermore, under the hypothesis that each class contains at least one cycle (if $n\geq 9$), then an upper bound heuristic is
$\psi_{c}(n)\leq k\approx\frac{n^{\frac{1}{2}}(n-1)}{2}=\frac{1}{2}n^{\frac{3}{2}}-O(n^{\frac{1}{2}}).$
@@
Let $n\geq 8$ and $c\colon E(K_{n})\rightarrow\{1,...,k\}$ a connected and complete edge-coloring of $K_{n}$ with $k=\psi_{c}(n)$ colors.
Let $x$ be the cardinality of the smallest chromatic class of $\varsigma$, that is, let $x=min\{\left|c^{-1}(i)\right|:i\in\left[k\right]\}$. Without loss of generality, suppose that $x=\left|c^{-1}(k)\right|$.
On the one hand, $c$ defines a partition of $E(K_{n})$ and it follows that $k\leq f_{n}(x):=n(n-1)/(2x)$.
On the other hand, $c$ is connected, therefore, $c^{-1}(k)$ is at least a tree subgraph. The maximum number of spanning trees in the subgraph $H=K_{x+1}$, such that $V(H)=V(c^{-1}(k))$, is at most $\frac{\binom{x+1}{2}}{x}$, then the chromatic class $x$ is incident to at most $\frac{\binom{x+1}{2}}{x}-1=\frac{x-1}{2}$ other chromatic classes. In addition, there are $(x+1)(n-(x+1))$ edges that are incident to some vertex of $c^{-1}(k)$ at least once, but on average each class is $\frac{2x}{x+1}$ incident to $c^{-1}(k)$. Since $c$ is complete, the number of chromatic classes incident to $c^{-1}(k)$ containing no edge in $H$ is at most
@@
Hence, there are at most $g_{n}(x)-1$ chromatic classes incident with some edge in $c^{-1}(k)$ where $g_{n}(x)-1:=\frac{(x+1)^{2}}{2x}(n-(x+1))+\frac{x-1}{2}$, according to the hypothesis of the average degree, i.e.
@@
and the first result is established.
For the second result, the hypothesis that each class contains at least one cycle implies that the average degree is at least 2. Equation 6.1 is modified as follows:
$\frac{(x+1)(n-x-1)}{2}$ (6.2)
Hence,
$g_{n}(x)=\frac{(n-x)(x+1)}{2}.$
In this case, equation $f_{n}(x_{0})=g_{n}(x_{0})$ is reduced to $n^{2}-n-x_{0}^{2}n-x_{0}n+x_{0}^{3}+x_{0}^{2}=0$. For the positive solution, we obtain $n=x_{0}^{2}$ (since $x_{0}\geq 3$, then $n\geq 9$) and then $x_{0}=\sqrt{n}$ taking the positive solution.
Since $f_{n}(x_{0})=\frac{n(n-1)}{2x_{0}}$, we get $f_{n}(x_{0})=\frac{n(n-1)}{2\sqrt{n}},$ and the last result is established. ∎
@@
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper $1$ $3$ $4$ $6$ $7$ $10$ $14$ $18$ $22$ $25$ Approx Upper $12$ $15$ $17$ $20$ Lower Improved $13$ $16$ $18$ Upper Heuristic $12$ $14$ $16$ Lower $1$ $3$ $4$ $6$ $7$ $10$ $11$ $12$ $13$ $14$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Upper $28$ $31$ $34$ $37$ $40$ $45$ $51$ $57$ $63$ $70$ Approx Upper $23$ $26$ $28$ $32$ $35$ $38$ $41$ $45$ $48$ $52$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Upper Heuristic $19$ $21$ $24$ $27$ $30$ $32$ $36$ $39$ $42$ $45$ Lower $19$ $26$ $27$ $28$ $29$ $30$ $31$ $32$ $33$ $42$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Upper $74$ $78$ $82$ $86$ $90$ $94$ $98$ $102$ $106$ $116$ Approx Upper $56$ $59$ $63$ $67$ $71$ $75$ $80$ $84$ $88$ $93$ Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Upper Heuristic $49$ $52$ $56$ $60$ $63$ $67$ $71$ $75$ $79$ $83$ Lower $43$ $44$ $45$ $46$ $47$ $48$ $49$ $50$ $51$ $93$
@@ 7 Conclusions
The problem at hand involves improving, by adapting a genetic algorithm called Rank GA, the known bounds of the connected pseudoachromatic index, denoted $\psi_{c}^{\prime}(K_{n})$, which represents the maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$. The Rank GA demonstrated significant effectiveness in improving $\psi_{c}^{\prime}(K_{n})$ bounds for complete graphs with connected classes. The algorithm was able to find the solutions due to its balance in exploring the solution space and exploiting promising solutions. The improved bounds of $\psi_{c}^{\prime}(K_{n})$ obtained by the Rank GA were validated against an analytical approximation, confirming their validity. The success of the Rank GA in enhancing the bounds of $\psi_{c}^{\prime}(K_{n})$ may significantly contribute to theoretical advancements in Chromatic Graph Theory. Upon examination of the resulting colorings, it is observed that the sizes of the chromatic classes vary, with most sizes equal to the minimum size stipulated by Theorem 6.1. However, their distribution does not align with the color patterns described in Theorem 6.2, meaning that they lack a projective plane structure in their arrangement.
@@
Appendix A Example Solutions
Figure 5 presents a sample solution with $n=12$ and 21 colors. Additional examples and a Python-based visualizer can be downloaded from ***https://www.dropbox.com/scl/fo/ibudsyzgwj0gz8gzs8ifi/AEXSShINNc23KMsURaxS9w4?rlkey=jsgye9loui8gapqnbm89c0csq&dl=0. The visualizer allows users to toggle the visualization of each color class to analyze solutions.
Figure 5: Example Solution

### v3

@@ Abstract
In this study, the Rank Genetic Algorithm was adapted to address a problem in the field of Chromatic Graph Theory, namely, on the parameter called the connected-pseudoachromatic index. We successfully improved several previously known bounds of that index of the complete graph.
Keywords: Genetic algorithm, rank GA, edge-coloring, connected classes, Hadwiger number.
@@ 1 Introduction
In recent decades, GAs have emerged as effective tools to address problems of graph theory with practical applications. Specifically, various GAs have been developed to tackle the Graph Coloring Problem (GCP), a fundamental issue in Graph Theory. In the simplest instance of the GCP, the objective is to assign colors to the vertices of a graph in such a way that no two adjacent vertices share the same color. The goal is to find a valid coloring of the graph with the minimal number of colors. Given its NP-hard nature and numerous practical applications, the GCP stands out as the most extensively researched combinatorial optimization problem in both computer science and mathematics [8]. In [12], we can find the most significant work in which GAs have been used to address the GCP. In addition to GCP, GAs have also been applied to other graph problems, among the most recent works are: the Shortest Path Problem [7], the Traveling Salesman Problem [14], Maximum Matching in bipartite graphs [11], and the Maximum Edge-Disjoint Path Problem [10].
@@
We are aware of the following studies where a genetic algorithm played a crucial role in addressing theoretical challenges: In [5] the Rank GA was used to find rainbow $t$-colorings of the family of Moore cages with girth 6: $(k;6)$, achieving an improvement in the upper bound for the 4-rainbow coloring of a $(4;6)$-cage. In addition, in [9], the Rank GA was adapted with recombination and mutation operators from graph theory. These operators enabled the discovery of Hajós operations to construct the symmetric 5-cycle from the complete symmetric digraph of order 3 using only Hajós operations. Finally, in [4] real solutions were identified by a genetic algorithm, thus contributing to the validation of a conjecture.
Our specific attention is directed towards the connected-pseudoachromatic index of the complete graph, denoted $\psi_{c}^{\prime}(K_{n})$ (for short $\psi_{c}(n)$). Determining $\psi_{c}(n)$ is difficult, and so far, there have been very few exact results for $\psi_{c}(n)$ of complete graphs with connected classes. However, it has been possible to bound it using analytical methods [2].
In this interdisciplinary study, our objective was to refine the known bounds of $\psi_{c}(n)$ exploiting the capabilities of the Rank GA. We successfully improved several of the previously known bounds. This synergy between Genetic Algorithms and mathematical theory contributes to the intersection of practical problem-solving and theoretical advancements.
The structure of this paper is as follows. In Section 2, we state the problem and include the necessary definitions to understand it. In Section 3, we describe the Rank GA and in Section 4, we present the obtained results. Finally, Section 5 contains the conclusions.
@@ 2 Problem statement
The maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$ that can be achieved is called the connected-pseudoachromatic index and is denoted $\psi_{c}(n)$. There are certain values of $n$ for which it is known, and for others, only the lower and upper bounds are known [1, 2]. The problem we address here, in particular, consists in improving the known bounds.
@@ 2.1 Edge-colorings of a complete graph
The connected-pseudoachromatic index $\psi^{\prime}_{c}(G)$ of a graph $G$ is the maximum number of colors $k$ for which a connected and complete edge-coloring exists in the graph $G$ using $k$ colors, when $G$ is connected; otherwise, it is defined as the maximum index over its connected components. When we color vertices and each chromatic class is connected, the parameter is also known as the Hadwiger number of $G$ from the perspective of minor graphs.
@@
In [2] it is shown that $\psi_{c}(n)$ can be bounded by analytical methods obtaining that $\psi_{c}(n)=\Theta(n^{3/2})$. Table 1 shows the known bounds of $\psi_{c}(n)$, for complete and connected graphs of $2\leq n\leq 31$ vertices. With the use of the computer, complete colorings can be achieved, and thus improve the lower bound of $\psi_{c}(n)$.
@@ 3.1 The Genetic Operators of the Rank GA
Once individuals are ranked based on their fitness, from best to worst, the Rank GA applies the following operations iteratively until a stopping criterion is met: Rank-based selection, Sort and rank, Rank-based recombination, Evaluate, sort, and rank, Rank-based mutation and Evaluate, sort, and rank.
@@ 3.1.3 Mutation by Rank
3.2 Pseudocode
The pseudocode for the Rank Genetic Algorithm (Rank GA), outlining the key steps involved in its execution, is shown in Algorithm 1 on page 1.
Algorithm 1 Rank Genetic Algorithm (Rank GA)
STEP 1: Initialization
Initialize population $\mathcal{P}$ randomly with $N$ individuals
Evaluate the fitness of each individual in $\mathcal{P}$
while stopping criterion not met do
     STEP 2: Rank-Based Selection
     Sort $\mathcal{P}$ by fitness in descending order
     Compute $r_{i}$ for each individual $i$ using equation (3.1)
     STEP 2.A: Cloning Based on Integer Part of Clone Number
     Initialize an empty list $\mathcal{P}^{\prime}$ to store clones
     for each individual $i$ in $\mathcal{P}$ do
         Calculate cloneNbri using equation (3.2)
         Generate floor(cloneNbri) clones of individual $i$ and add them to $\mathcal{P}^{\prime}$
     end for
     STEP 2.B: Fractional Cloning (Cyclic)
     Initialize $i=0$
     while size of $\mathcal{P}^{\prime}$ is less than $N$ do
         Let $f_{i}$ be the fractional part of cloneNbri
         if random number $<$ $f_{i}$ then
              Generate one additional clone of individual $i$ and add it to $\mathcal{P}^{\prime}$
         end if
         Increment $i$ by 1 (modulo $N$) to cycle through the population
     end while
     Replace original population $\mathcal{P}$ with the cloned population $\mathcal{P}^{\prime}$
     STEP 3: Rank-Based Recombination
     Sort $\mathcal{P}$ by fitness in descending order
     Compute $r_{i}$ for each individual $i$ using equation (3.1)
     for each pair of individuals $i$ and $i+1$ in $\mathcal{P}$ do
         Perform recombination between individuals $i$ and $i+1$
         Replace the parents $i$ and $i+1$ with the two offspring
     end for
     Evaluate the fitness of each individual in $\mathcal{P}$
     STEP 4: Mutation by Rank
     Sort $\mathcal{P}$ by fitness in descending order
     Compute $r_{i}$ for each individual $i$ using equation (3.1)
     for each individual $i$ in $\mathcal{P}$ do
         Calculate mutation probability $p_{i}$ using equation (3.3)
         for each gene $g$ in the genotype of individual $i$ do
              if random number $<$ $p_{i}$ then
                  Mutate gene $g$
              end if
         end for
     end for
     Evaluate the fitness of each individual in $\mathcal{P}$
end while
STEP 5: Return the Best Individual
Return the best individual in $\mathcal{P}$
@@ 4 Results
4.1 On the lower bound
Table 2 presents the improved lower bounds obtained through the Rank GA for various values of $n$. The known upper bounds for $\psi_{c}(n)$ are also shown for reference in Table 2. By comparing the improved lower bounds with the existing lower and upper bounds, we observed significant reductions in the gap between them.
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper 1 3 4 6 7 10 $14$ $18$ $22$ $25$ Lower Improved $13$ $16$ $18$ Lower 1 3 4 6 7 10 $11$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Upper $28$ $31$ $34$ $37$ $40$ $45$ $51$ $57$ $63$ $70$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Lower $26$ $27$ $28$ $29$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Upper $74$ $78$ $82$ $86$ $90$ $94$ $98$ $102$ $106$ $116$ Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Lower $93$
Table 2: Improved lower bounds with the Rank GA for $\psi_{c}(n)$ and $n\leq 31$.
@@ 4.2 On the upper bound
Let $n\geq 8$ and $\varsigma\colon E(K_{n})\rightarrow\{1,\dots,k\}$ a connected and complete edge-coloring of $K_{n}$ with $k=\psi_{c}(n)$ colors.
Let $x$ be the cardinality of the smallest chromatic class of $\varsigma$, that is, let $x=min\{\left|\varsigma^{-1}(i)\right|:i\in\{1,\dots,k\}\}$. Without loss of generality, suppose that $x=\left|\varsigma^{-1}(k)\right|$.
On the one hand, $\varsigma$ defines a partition of $E(K_{n})$ and it follows that $k\leq f_{n}(x):=n(n-1)/(2x)$.
On the other hand, $\varsigma$ is connected, therefore, $\varsigma^{-1}(k)$ is at least a tree subgraph. The maximum number of spanning trees in the subgraph $H=K_{x+1}$, such that $V(H)=V(\varsigma^{-1}(k))$, is at most $\frac{\binom{x+1}{2}}{x}$, then the chromatic class $x$ is incident to at most $\frac{\binom{x+1}{2}}{x}-1=\frac{x-1}{2}$ other chromatic classes. In addition, there are $(x+1)(n-(x+1))$ edges that are incident to some vertex of $\varsigma^{-1}(k)$ at least once, but on average each class is $\frac{2x}{x+1}$ incident to $\varsigma^{-1}(k)$. Since $\varsigma$ is complete, the number of chromatic classes incident to $\varsigma^{-1}(k)$ containing no edge in $H$ is at most
@@
Hence, there are at most $g_{n}(x)-1$ chromatic classes incident with some edge in $\varsigma^{-1}(k)$ where $g_{n}(x)-1:=\frac{(x+1)^{2}}{2x}(n-(x+1))+\frac{x-1}{2}$, according to the hypothesis of the average degree, i.e.
@@
and the result is established. ∎
@@
$n$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $11$ Upper 1 3 4 6 7 10 Approx Upper $12$ $15$ $17$ $20$ Lower Improved $13$ $16$ $18$ Lower 1 3 4 6 7 10 $11$ $n$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$ $21$ Approx Upper $23$ 26 $28$ $32$ $35$ $38$ $41$ $45$ $48$ $52$ Lower Improved $21$ $32$ $35$ $37$ $39$ $43$ Lower 26 $27$ $28$ $29$ $n$ $22$ $23$ $24$ $25$ $26$ $27$ $28$ $29$ $30$ $31$ Approx Upper $56$ $59$ $63$ $67$ $71$ $75$ $80$ $84$ $88$ 93 Lower Improved $46$ $50$ $52$ $54$ $55$ $59$ $62$ $64$ $66$ Lower 93
@@ 5 Conclusions
The problem at hand involves improving, by adapting a genetic algorithm called Rank GA, the known bounds of the connected-pseudoachromatic index for the complete graph of order $n$, denoted by $\psi_{c}(n)$, which represents the maximum number of colors in a connected and complete coloring of a complete graph $K_{n}$ of order $n$. The Rank GA demonstrated significant effectiveness in improving $\psi_{c}(n)$ bounds for complete graphs with connected classes. Our paper enters into the set of problems with results for small $n$ values such that: [13] where it was proved that the well-known Erdő-Faber-Lóvasz conjecture is true for $n\leq 12$; [15] where the extremal values of the $C_{4}$-free graphs are shown for $n\leq 31$. While in [6] all the Steiner Triplets Systems of order 19 were found.
The algorithm was able to find the solutions due to its balance in exploring the solution space and exploiting promising solutions. The improved bounds of $\psi_{c}(n)$ obtained by the Rank GA were validated against an analytical approximation, confirming their validity. The success of the Rank GA in enhancing the bounds of $\psi_{c}(n)$ may significantly contribute to theoretical advancements in Chromatic Graph Theory. Upon examination of the resulting colorings, it is observed that the sizes of the chromatic classes vary, with most sizes equal to the minimum size stipulated by Theorem 4.1. However, their distribution does not align with the color patterns described in Theorem 4.2, meaning that they lack a projective plane structure in their arrangement.
@@
Finally, Figure 5 presents a sample solution with $n=12$ and 21 colors. Additional examples and a Python-based visualizer can be downloaded from ***https://www.dropbox.com/scl/fo/ibudsyzgwj0gz8gzs8ifi/AEXSShINNc23KMsURaxS9w4?rlkey=jsgye9loui8gapqnbm89c0csq&dl=0. The visualizer allows users to toggle the visualization of each color class to analyze solutions.
Figure 5: Example solution.